# Many-Body Localization in the Age of Classical Computing

### v2

@@ VII.3.1 Proximity to Anderson Insulator
Whether the Anderson localization may persist in the presence of interaction was the original formulation of the MBL problem [51]. The possibility of investigating MBL in the disordered XXZ spin chain (7) from the perspective of this model’s non-interacting ($\Delta=0$) limit allows for non-trivial insights into the physics of strongly disordered spin chains.
Even in the ergodic regime of disordered XXZ spin chain, e.g., at $W=1.5$ and $L=16$, the spectral function $f^{2}(\omega)$ (defined in Eq. 41) of the non-interacting model ($\Delta=0$) provides, at sufficiently large frequencies $\omega$, a good approximation of $f^{2}(\omega)$ in the presence of interactions ($\Delta=1$). This demonstrates the resemblance of the dynamics on relatively short time scales in the interacting and non-interacting cases. Significant differences between $\Delta=0$ and $\Delta=1$ appear only at small frequencies $\omega$, corresponding to evolution on longer time scales, c.f. Fig. 10(a). The presence of the Anderson localized phase at $\Delta=0$ allowed [678] to propose a phenomenological model describing the behavior of the spectral function $f^{2}(\omega)$ at small frequencies.
The main premise is that the Anderson LIOMs $\hat{Q}_{j}$, corresponding to the localized single-particle orbitals of the Anderson insulator, may determine the memory of the initial state kept in the system at late times, and hence give rise to anomalously slow dynamics. In the strict non-interacting case, $\Delta=0$, the Anderson LIOMs contribute to the $\omega\to 0$ limit of the spectral function as $f^{2}(\omega)\stackrel{{\scriptstyle\omega\ll 1}}{{=}}C_{0}\delta(\omega)$. When the interactions are introduced, $\Delta=1$, the Anderson LIOMs are no longer stable and decay in time. Calculation of the correlator $\braket{\hat{Q}_{j}(t)\hat{Q}_{j}}$ allows to extract a time scale $\tau$ at which the Anderson LIOM $Q_{j}$ decays. The key finding of [678] is that the distribution $f(\tau)$ of the time scales $\tau$ becomes extremely broad with increasing $W$, and is well captured, in a wide range of $\tau$, by a power-law dependence $f(\tau)\propto\tau^{-\mu}$, with an exponent $\mu$ dependent on the disorder strength $W$. The finite lifetime of Anderson LIOMs amounts to broadening the Delta peaks in $f^{2}(\omega)$ at $\omega=0$ present in the non-interacting case. Assuming a Lorentzian broadening of the Delta peaks, the following low-frequency behavior of the spectral function was obtained [678]
@@
where $\tau_{\mathrm{min}}$ and $\tau_{\mathrm{max}}$ correspond to minimal and maximal time scales at which the weight of the distribution $f(\tau)$ becomes negligible. For $\omega\ll\tau_{\mathrm{max}}^{-1}$, the spectral function has a plateau, $f^{2}(\omega)\propto\mathrm{const}$, which can be used to define the Thouless energy, see Sec. VI.1. The plateau appears in the ergodic regime, $W<W^{T}(L)$, and shrinks with the growth of $W$ as $\tau_{\mathrm{max}}$ increases. When $\tau_{\mathrm{max}}$ approaches the Heisenberg time $t_{H}$ at $W\approx W^{T}(L)$ (for instance, in disordered XXZ spin chain at $L=16$ and $W\approx 2$) the plateau disappears. At that point, the exponent $\mu$ is close to unity, which corresponds to the $f^{2}(\omega)\propto 1/\omega$ behavior observed in [416, 572] for the imbalance or the local $\hat{S}_{j}^{z}$ operator. The exponent $\mu$ increases upon a further increase of $W$, corresponding to behavior $f^{2}(\omega)\propto 1/\omega^{2-\mu}$ observed in Fig. 8(b), and the regime of the slow persistent decays of local autocorrelation functions emerges.
The main consequence of the phenomenological model of the spectral function (52) is that the dynamics of disordered spin chains may be perceived as being governed, in a wide interval of $W$, by the broad distribution of the relaxation times of Anderson LIOMs. Even in the MBL regime, $W>W^{*}(L)$, $\tau_{\mathrm{min}}$ is significantly smaller than $t_{H}$, while $\tau_{\mathrm{max}}$ exceeds the Heisenberg time. Accordingly, in that regime, a certain fraction of Anderson LIOMs remain stable and the remaining fraction delocalize, corresponding to the non-trivial time dynamics observed in the MBL regime, see, e.g., Fig. 11(a).
@@
$\hat{H}_{0}=\sum_{\alpha=1}^{L}\epsilon_{\alpha}Q_{\alpha}$ (53)
is diagonal in the basis of Anderson orbitals, and the interaction part $H_{\mathrm{int}}\equiv\sum_{j=1}^{L}\Delta S^{z}_{j}S^{z}_{j+1}$ is a perturbation of the Anderson model [^29] We note that we express the XXZ Hamiltonian in terms of spin-1/2 operators, while the notion of single-particle Anderson orbitals is usually invoked in therms of spinless fermion Hamiltonians, such as the one in Eq. (5), onto which the spin-1/2 Hamiltonians can be mapped via the Jordan-Wigner transformation.. The important observation of [341] is that the interaction term can be split into two orthogonal parts as
@@
where the orthogonality is defined by the Hilbert-Schmidt product, $\braket{A,B}=\frac{1}{\mathcal{N}}\mathrm{tr}\left(A^{{\dagger}}B\right)$, and ${\cal N}$ is the corresponding Hilbert space dimension. The division of the interaction term is performed in such a way that $[\hat{H}^{\parallel}_{\mathrm{int}},Q_{j}]=0$, i.e., the Anderson LIOMs commute with $H^{\parallel}_{\mathrm{int}}$. Consequently, only the term $\hat{H}^{\perp}_{\mathrm{int}}$ is a genuine non-trivial perturbation to the Anderson insulator.
In the limit of large disorder strength $W$, the Anderson LIOMs $Q_{j}$ become identical to the local spin operators $S^{z}_{j}$. In that limit, the interaction term $\hat{H}_{\mathrm{int}}=\sum_{j}\Delta S^{z}_{j}S^{z}_{j+1}$ already commutes with Anderson LIOMs, and $\hat{H}^{\perp}_{\mathrm{int}}=0$ for $W\to\infty$. The Anderson LIOMs get broadened for finite disorder strengths, and the perturbation to Anderson insulator, $\hat{H}^{\parallel}_{\mathrm{int}}$, is non-vanishing. Nevertheless, at strong disorder $W$, the interaction term, $\Delta S^{z}_{j}S^{z}_{j+1}$, is still largely parallel to the Anderson LIOMs due to their exponential localization. Quantitatively, Ref. [341] shows that $||H^{\parallel}_{\mathrm{int}}||\propto 1/W$ (where $||A||=\sqrt{\braket{A,A}}$), i.e., the norm of the true perturbation of the Anderson insulator decays as $1/W$. In contrast, the norm of the non-interacting Hamiltonian increases as $||\hat{H}_{0}||\propto W$.
Therefore, at sufficiently large $W$, the perturbation $\hat{H}^{\perp}_{\mathrm{int}}$ becomes rather weak, and the disordered XXZ spin chain may be viewed as a weakly perturbed Anderson insulator. A weak but non-vanishing perturbation results in slow dynamics towards the thermal equilibrium, sharing certain similarities with the weakly perturbed integrable models [417, 630], posing a challenge to understanding the fate of disordered spin chains in the $L,t\to\infty$ limit. Instead, when the perturbation term $\hat{H}^{\perp}_{\mathrm{int}}$ is rescaled to match the norm of $\hat{H}_{0}$, the ensuing model remains ergodic at any disorder strength $W$ [341]. Similarly, the ergodicity is restored upon an introduction of a strong pair hopping term [342], which introduces non-trivial perturbation to the Anderson insulator with norm $||\hat{H}^{\perp}_{\mathrm{int}}||$ of the similar value as the norm of $\hat{H}_{0}$.
The discussed results outline the perspective on disordered spin chains provided by the presence of the non-interacting, Anderson-localized limit of the disordered spin chains. Further investigations of the interplay of Anderson LIOMs $Q_{\alpha}$ with interactions that induce their gradual decay may lead to new insights into the physics of slow dynamics in strongly disordered spin chains.
@@ VII.3.2 Stretched exponential decays
The power-law decay of local correlators may be attributed to rare fluctuations in the disorder, i.e., to Griffiths regions, which acts as bottlenecks for transport [7]. The observed stretched exponential decay at $W\approx W^{T}(L)$ may be explained by resonances between many-body states with locally differing magnetization patterns [146]. Such resonances manifest as large oscillations of local correlators and average out when the disorder average is taken, c.f. Fig. 10. This motivated [380] to propose, without making any predictions about the value of $W_{C}$ nor about the mechanism of the possible transition to MBL phase, that many-body resonances characterize thermalization in most of the prethermal MBL regime, i.e., at disorder strengths $W^{T}(L)\lessapprox W\lessapprox W_{C}$ at which the system ultimately thermalizes in the $L,t\to\infty$ limit. While the rare-region effects may plan an important role at larger disorder strengths,  [380] suggested that the rare-region effects are not relevant at those intermediate disorder strengths. The argument relies on the absence of the fat power-law tails in the resistance distribution [fat tails are consistent with the presence of Griffiths rare regions in the disordered XXZ spin chain [568]], and on the fact that power-law decays are also observed in quasiperiodic systems [173] in which the rare-regions are absent, see also Sec. X.2. Importantly, the many-body resonances are proposed as an intermediate step towards thermalization associated with establishing a local equilibrium. At longer times, the hydrodynamic behavior characterized by the power-law decays of local observables is expected to take over.
@@
The claim about stretched exponential decay is substantiated numerically by the Jacobi algorithm for iterative matrix diagonalization [236]. The algorithm, see also [381], identifies the largest off-diagonal matrix element of the currently considered rotated Hamiltonian $H(j)$. It performs a unitary transformation $R$ to eliminate the identified element from the rotated matrix, $H(j+1)=R^{{\dagger}}H(j)R$, so that the off-diagonal weight of $H(j)$ decreases at each step $j$. Resonances between respective many-body states are found when the decimated off-diagonal matrix element is much larger than the corresponding difference in diagonal elements. This observation allows [380] to extract a broad, power-law distribution of resonance frequencies and argue that the hierarchy of successive resonances leads to the stretched exponential decay of autocorrelators, unless the system flows towards the MBL phase and the dynamics become arrested.
@@
Nevertheless, the stretched exponential prediction breaks down, for $L=20$, already at $W=3$ for the disordered XXZ spin chain. Indeed, as shown in Fig. 10(b), the decay of $C(t)$ may be well fitted by a power-law or logarithmic dependence in time [^30] The decay of $C(t)$ at $W=3$ is consistent, for instance, with a power-law. The amount by which $C(t)$ decays in the available time interval precludes any strong statements about the functional form of the decay.. Still, a comparison of the $W=3$ case with the behavior of $C(t)$ at $W=2$, which is well fitted by a stretched exponential [380], indicates a possibility for the stretched exponential decay of $C(t)$ to be observed at $W=3$ for system size slightly larger than $L=20$. An implication of the assumed stretched exponential decay for $W=8$ would be that the slow decay of $C(t)$, well fitted by a power-law in Fig. 11(a), could become a faster decay at longer times. However, even under that assumption, the time scale of thermalization for a stretched decay consistent with the results in Fig. 11(a) would be enormous, and the system would remain non-ergodic for any practical purpose. The broad distribution of time scales associated with the hierarchy of resonances uncovered by the Jacobi rotations is a property of strongly disordered spin chains similar to the broad distribution of the decay times of Anderson LIOMs [678]. Understanding the relation between the two approaches is an interesting direction for further investigations of the regime of slow dynamics in strongly disordered spin chains.
@@ VII.3.3 Internal clock of many-body dynamics
Searching for a paradigm to provide a universal description of the regime of slow dynamics in the disordered spin chains, Ref. [198] makes the following observation. Considering a fixed value of disorder strength $W$, the dynamics of a specified observable $A(t)$ for two particular disorder realizations differ due to sample-to-sample fluctuations in the properties of the system. Extraction of universal trends in the dynamics of the observable $A$ could be more straightforward if one introduces an internal clock that measures the flow of a fictitious time $\Gamma$. If the quantity $\Gamma$ is appropriately chosen, the dynamics of $A$, expressed as a function of $\Gamma$, should show at least a partial synchronization between the samples, simplifying the time evolution analysis.
As the internal clock, [198] proposed to employ the entanglement entropy
@@
Deep in the ergodic regime of disordered spin chains, the entanglement spreads ballistically, $S(t)\propto t$, and the fictitious time measured by the internal clock is proportional to the real-time, $\Gamma\propto t$. With the increase in the disorder strength, the relation between the entanglement entropy and time ceases to be linear. Additionally, the fictitious time $\Gamma$ flows differently for separate disorder realizations, and the relation between the considered observable $A$ and $\Gamma$ becomes non-trivial. Focusing on the density autocorrelation function $C(t)$, and its fluctuation $\mathcal{F}$ in a time interval of a fixed length $\Delta t$, Ref. [198] demonstrates that the sample-to-sample fluctuations of $C(\Gamma)$ and $\mathcal{F}(\Gamma)$ are significantly reduced compared to $C(t)$ and $\mathcal{F}(t)$.
The main finding of [198] is that the disorder averages of $C(t)$ and $\mathcal{F}(t)$ plotted as functions of the disorder-averaged entanglement entropy $S(t)$ collapse onto a single master curve which depends only on the system size for different values of $W<8$ in the disordered XXZ spin chain. Thus, $S(t)$ works well as the internal clock $\Gamma$, enabling the identification of a high degree of universality in the regime of slow dynamics in the disordered XXZ spin chain. Consequently, in the considered interval of system sizes and time scales, there are no visible signatures of the MBL phase. Independently of the disorder strength $W$, the autocorrelation function $C(t)$ decay is dictated by the flow of the fictitious time $\Gamma$, i.e., the increase of entanglement entropy $S(t)$. In contrast, an MBL phase would require the emergence of another branch of the master curve for $L\to\infty$, along which the fictitious time grows indefinitely (since $S(t)\propto\log(t)$), and the autocorrelation functions do not decay. The absence of such a branch of the master curve at the time scale $t<10^{3}$ considered in [198] is consistent with the persistent decay of the autocorrelation function $C(t)$ shown in Fig. 11(a). Therefore, a smoking gun evidence for the MBL phase, if it exists, should be looked for at larger time scales and system sizes.

### v3

@@ VII.3.1 Proximity to Anderson Insulator
Whether the Anderson localization may persist in the presence of interactions was the original formulation of the MBL problem [54]. The possibility of investigating MBL in the disordered XXZ spin chain (7) from the perspective of this model’s non-interacting limit ($\Delta=0$) allows for non-trivial insights into the physics of strongly disordered spin chains.
Even in the ergodic regime of the disordered XXZ spin chain, e.g., at $W=1.5$ and $L=16$, the spectral function $f^{2}(\omega)$ (defined in Eq. (41)) of the non-interacting model ($\Delta=0$) provides, at sufficiently large frequencies $\omega$, a good approximation of $f^{2}(\omega)$ in the presence of interactions ($\Delta=1$). This demonstrates the resemblance of the dynamics on relatively short time scales in the interacting and non-interacting cases. Significant differences between $\Delta=0$ and $\Delta=1$ appear only at small frequencies $\omega$, corresponding to evolution on longer time scales, c.f. Fig. 10(a). The presence of the Anderson localized phase at $\Delta=0$ allowed [733] to propose a phenomenological model describing the behavior of the spectral function $f^{2}(\omega)$ at small frequencies.
The main premise is that the Anderson LIOMs $\hat{Q}_{j}$, corresponding to the localized single-particle orbitals of the Anderson insulator, may determine the memory of the initial state kept in the system at late times, and hence give rise to anomalously slow dynamics. In the strict non-interacting case, $\Delta=0$, the Anderson LIOMs contribute to the $\omega\to 0$ limit of the spectral function as $f^{2}(\omega)\stackrel{{\scriptstyle\omega\ll 1}}{{=}}C_{0}\delta(\omega)$. When the interactions are introduced, $\Delta=1$, the Anderson LIOMs are no longer stable and decay in time. Calculation of the correlator $\braket{\hat{Q}_{j}(t)\hat{Q}_{j}}$ allows to extract a time scale $\tau$ at which the Anderson LIOM $\hat{Q}_{j}$ decays. The key finding of [733] is that the distribution $w(\tau)$ of the numerically extracted relaxation time scales $\tau$ becomes extremely broad with increasing $W$, and is well captured, in a wide range of $\tau$, by a power-law dependence $w(\tau)\propto\tau^{-\mu}$, with an exponent $\mu$ dependent on the disorder strength $W$. The finite lifetime of Anderson LIOMs amounts to the broadening of the Delta peaks in $f^{2}(\omega)$ at $\omega=0$ present in the non-interacting case. Assuming a Lorentzian broadening of the Delta peaks, the following low-frequency behavior of the spectral function was obtained [733],
@@
where $\tau_{\mathrm{min}}$ and $\tau_{\mathrm{max}}$ correspond to minimal and maximal time scales at which the weight of the distribution $w(\tau)$ becomes negligible. For $\omega\ll\tau_{\mathrm{max}}^{-1}$, the spectral function has a plateau, $f^{2}(\omega)\propto\mathrm{const}$, which can be used to define the Thouless energy, see Sec. VI.1. The plateau appears in the ergodic regime, $W<W^{T}(L)$, and shrinks with the growth of $W$ as $\tau_{\mathrm{max}}$ increases. When $\tau_{\mathrm{max}}$ approaches the Heisenberg time $t_{H}$ at $W\approx W^{T}(L)$ (for instance, in the disordered XXZ spin chain at $L=16$ and $W\approx 2$), the plateau disappears. At that point, the exponent $\mu$ is close to unity, which corresponds to the $f^{2}(\omega)\propto 1/\omega$ behavior observed in [452, 619] for the imbalance or the local $\hat{S}_{j}^{z}$ operator. The exponent $\mu$ increases upon a further increase of $W$, corresponding to the behavior $f^{2}(\omega)\propto 1/\omega^{2-\mu}$ observed in Fig. 8(b), and the regime of the slow persistent decays of local autocorrelation functions emerges.
The main consequence of the phenomenological model of the spectral function (52) is that the dynamics of disordered spin chains may be perceived as being governed, in a wide interval of $W$, by the broad distribution of the relaxation times of Anderson LIOMs. Even in the MBL regime, $W>W^{*}(L)$, $\tau_{\mathrm{min}}$ is significantly smaller than $t_{H}$, while $\tau_{\mathrm{max}}$ exceeds the Heisenberg time. Accordingly, in that regime, a certain fraction of Anderson LIOMs remain stable and the remaining fraction delocalize, corresponding to the non-trivial dynamics observed in the MBL regime, see, e.g., Fig. 11(a).
@@
$\hat{H}_{0}=\sum_{\alpha=1}^{L}\epsilon_{\alpha}\hat{Q}_{\alpha}$ (53)
is diagonal in the basis of Anderson orbitals, and the interaction part $H_{\mathrm{int}}\equiv\sum_{j=1}^{L}\Delta\hat{S}^{z}_{j}\hat{S}^{z}_{j+1}$ is a perturbation of the Anderson model [^38] We note that we express the XXZ Hamiltonian in terms of spin-1/2 operators, while the notion of single-particle Anderson orbitals is usually invoked in therms of spinless fermion Hamiltonians, such as the one in Eq. (5), onto which the spin-1/2 Hamiltonians can be mapped via the Jordan-Wigner transformation.. The important observation of [370] is that the interaction term can be split into two orthogonal parts as
@@
where the orthogonality is defined by the Hilbert-Schmidt product, $\braket{A,B}=\frac{1}{\mathcal{N}}\mathrm{tr}\left(A^{{\dagger}}B\right)$, and ${\cal N}$ is the corresponding Hilbert space dimension. The division of the interaction term is performed in such a way that $[\hat{H}^{\parallel}_{\mathrm{int}},\hat{Q}_{j}]=0$, i.e., the Anderson LIOMs commute with $\hat{H}^{\parallel}_{\mathrm{int}}$. Consequently, only the term $\hat{H}^{\perp}_{\mathrm{int}}$ is a genuine non-trivial perturbation to the Anderson insulator.
In the limit of large disorder strength $W$, the Anderson LIOMs $\hat{Q}_{j}$ become identical to the local spin operators $\hat{S}^{z}_{j}$. In that limit, the interaction term $\hat{H}_{\mathrm{int}}=\sum_{j}\Delta\hat{S}^{z}_{j}\hat{S}^{z}_{j+1}$ already commutes with Anderson LIOMs, and $\hat{H}^{\perp}_{\mathrm{int}}=0$ for $W\to\infty$. The Anderson LIOMs get broadened for finite disorder strengths, and the perturbation to Anderson insulator, $\hat{H}^{\parallel}_{\mathrm{int}}$, is non-vanishing. Nevertheless, at strong disorder $W$, the interaction term, $\Delta\hat{S}^{z}_{j}\hat{S}^{z}_{j+1}$, is still largely parallel to the Anderson LIOMs due to their exponential localization. Quantitatively, [370] showed that $||\hat{H}^{\parallel}_{\mathrm{int}}||\propto 1/W$ (where $||A||=\sqrt{\braket{A,A}}$), i.e., the norm of the true perturbation of the Anderson insulator decays as $1/W$. In contrast, the norm of the non-interacting Hamiltonian increases as $||\hat{H}_{0}||\propto W$.
Therefore, at sufficiently large $W$, the perturbation $\hat{H}^{\perp}_{\mathrm{int}}$ becomes rather weak, and the disordered XXZ spin chain may be viewed as a weakly perturbed Anderson insulator. A weak but non-vanishing perturbation results in slow dynamics towards the thermal equilibrium, sharing certain similarities with the weakly perturbed integrable models [453, 680], posing a challenge to understanding the fate of the disordered spin chains in the $L,t\to\infty$ limit. Instead, when the perturbation term $\hat{H}^{\perp}_{\mathrm{int}}$ is rescaled to match the norm of $\hat{H}_{0}$, the ensuing model remains ergodic at any disorder strength $W$ [370]. Similarly, the ergodicity is restored upon an introduction of a strong pair hopping term [371], which introduces non-trivial perturbation to the Anderson insulator with norm $||\hat{H}^{\perp}_{\mathrm{int}}||$ of the similar value as the norm of $\hat{H}_{0}$.
The discussed results outline the perspective on the disordered spin chains provided by the presence of the non-interacting, Anderson-localized limit. Further investigations of the interplay of Anderson LIOMs $\hat{Q}_{\alpha}$ with interactions that induce their gradual decay may lead to new insights into the physics of slow dynamics in strongly disordered spin chains.
@@ VII.3.2 Stretched exponential decays
The power-law decay of local correlators may be attributed to rare fluctuations in the disorder, i.e., to Griffiths regions, which act as bottlenecks for transport [7]. The observed stretched exponential decay at $W\approx W^{T}(L)$ may be explained by the resonances between many-body states with locally differing magnetization patterns [151]. Such resonances manifest themselves as large oscillations of local correlators and average out when the disorder average is taken, c.f. Fig. 10. This motivated [416] to propose, without making any predictions about the value of $W_{C}$ nor about the mechanism of the possible transition to MBL phase, that many-body resonances characterize thermalization in most of the prethermal MBL regime, i.e., at disorder strengths $W^{T}(L)\lessapprox W\lessapprox W_{C}$ at which the system ultimately thermalizes in the $L,t\to\infty$ limit. In contrast, while the rare-region effects may play an important role at larger disorder strengths,  [416] suggested that the rare-region effects are not relevant at those intermediate disorder strengths. The argument relies on the absence of the fat power-law tails in the resistance distribution [fat tails are consistent with the presence of Griffiths rare regions in the disordered XXZ spin chain [615]], and on the fact that power-law decays are also observed in quasiperiodic systems [182] in which the rare-regions are absent, see also Sec. X.2. Importantly, the many-body resonances are proposed as an intermediate step towards thermalization associated with establishing a local equilibrium. At longer times, the hydrodynamic behavior characterized by the power-law decays of local observables is expected to take over.
@@
The claim about stretched exponential decay is substantiated numerically by the Jacobi algorithm for iterative matrix diagonalization [252]. The algorithm, see also [417], identifies the largest off-diagonal matrix element of the currently considered rotated Hamiltonian $\hat{H}(j)$. It performs a unitary transformation $\hat{R}$ to eliminate the identified element from the rotated matrix, $\hat{H}(j+1)=\hat{R}^{{\dagger}}\hat{H}(j)\hat{R}$, so that the off-diagonal weight of $\hat{H}(j)$ decreases at each step $j$. Resonances between respective many-body states are found when the decimated off-diagonal matrix element is much larger than the corresponding difference in diagonal elements. This observation allowed [416] to extract a broad, power-law distribution of resonance frequencies and argue that the hierarchy of successive resonances leads to the stretched exponential decay of autocorrelators, unless the system flows towards the MBL phase and the dynamics become arrested.
@@
Nevertheless, the stretched exponential prediction breaks down, for $L=20$, already at $W=3$ for the disordered XXZ spin chain. Indeed, as shown in Fig. 10(b), the decay of $C(t)$ may be well fitted by a power-law or logarithmic dependence in time [^39] The decay of $C(t)$ at $W=3$ is consistent, for instance, with a power-law. The amount by which $C(t)$ decays in the available time interval precludes any strong statements about the functional form of the decay.. Still, a comparison of the $W=3$ case with the behavior of $C(t)$ at $W=2$, which is well fitted by a stretched exponential [416], indicates a possibility for the stretched exponential decay of $C(t)$ to be observed at $W=3$ for system size slightly larger than $L=20$. An implication of the assumed stretched exponential decay for $W=8$ would be that the slow decay of $C(t)$, well fitted by a power-law in Fig. 11(a), could become a faster decay at longer times. However, even under that assumption, the time scale of thermalization for a stretched decay, consistent with the results in Fig. 11(a), would be enormous, and the system would remain non-ergodic for any practical purpose. The broad distribution of time scales associated with the hierarchy of resonances uncovered by the Jacobi rotations is a property of strongly disordered spin chains similar to the broad distribution of the decay times of Anderson LIOMs [733]. Understanding the relation between the two approaches is an interesting direction for further investigations of the regime of slow dynamics in strongly disordered spin chains.
@@ VII.3.3 Internal clock of many-body dynamics
Searching for a paradigm to provide a universal description of the regime of slow dynamics in the disordered spin chains, [208] made the following observation. Considering a fixed value of disorder strength $W$, the dynamics of a specified observable $A(t)$ for two particular disorder realizations differ due to sample-to-sample fluctuations in the properties of the system. Extraction of universal trends in the dynamics of the observable $A$ could be more straightforward if one introduces an internal clock that measures the flow of a fictitious time $\Gamma$. If the quantity $\Gamma$ is appropriately chosen, the dynamics of $A$, expressed as a function of $\Gamma$, should show at least a partial synchronization between the samples, simplifying the time evolution analysis.
As the internal clock, [208] proposed to employ the entanglement entropy,
@@
Deep in the ergodic regime of disordered spin chains, the entanglement spreads ballistically, $S(t)\propto t$, and the fictitious time measured by the internal clock is proportional to the real-time, $\Gamma\propto t$. With the increase in the disorder strength, the relation between the entanglement entropy and time ceases to be linear. Additionally, the fictitious time $\Gamma$ flows differently for separate disorder realizations, and the relation between the considered observable $A$ and $\Gamma$ becomes non-trivial. Focusing on the density autocorrelation function $C(t)$, and its fluctuation $\mathcal{F}$ in a time interval of a fixed length $\Delta t$, [208] demonstrated that the sample-to-sample fluctuations of $C(\Gamma)$ and $\mathcal{F}(\Gamma)$ are significantly reduced compared to $C(t)$ and $\mathcal{F}(t)$.
The main finding of [208] is that the disorder averages of $C(t)$ and $\mathcal{F}(t)$, plotted as functions of the disorder-averaged entanglement entropy $S(t)$, collapse onto a single master curve which depends only on the system size for different values of $W<8$ in the disordered XXZ spin chain. Thus, $S(t)$ works well as the internal clock $\Gamma$, enabling the identification of a high degree of universality in the regime of slow dynamics in the disordered XXZ spin chain. Consequently, in the considered interval of system sizes and time scales, there are no visible signatures of the MBL phase. Independently of the disorder strength $W$, the autocorrelation function $C(t)$ decay is dictated by the flow of the fictitious time $\Gamma$, i.e., the increase of entanglement entropy $S(t)$. In contrast, an MBL phase would require the emergence of another branch of the master curve for $L\to\infty$, along which the fictitious time grows indefinitely (since $S(t)\propto\log(t)$), and the autocorrelation functions do not decay. The absence of such a branch of the master curve at the time scale $t<10^{3}$ considered in [208] is consistent with the persistent decay of the autocorrelation function $C(t)$ shown in Fig. 11(a). Therefore, a smoking gun evidence for the MBL phase, if it exists, should be looked for at larger time scales and system sizes.